OpenStudy (anonymous):

The population of California was 10,586,223 in 1950 and 23,668,562 in 1980. Assume that the population grows exponentially. (a) Find a function that models the population t years after 1950. (Round rate r to four decimal places.) (b) Find the time required for the population to double. (Round your answer to the nearest whole number.) (c) Use the function from part (a) to predict the population of California in the year 2003.

OpenStudy (anonymous):

nice numbers...

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

easy way or hard way?

OpenStudy (anonymous):

Whatever you want to do. I just want the answers

OpenStudy (anonymous):

probably need the hard way , probably your teacher wants \[P(t)=P_0e^{rt}\]

OpenStudy (anonymous):

yeah that is what he wants

OpenStudy (anonymous):

ok first we do it the hard way. then if you want i can show you and easier way, or we can just be done. set \[23,668,562=10,586,223 e^{30r}\] and solve for r. is this step clear?

OpenStudy (anonymous):

that is an o and not a zero, correct? Yes that is clear

OpenStudy (anonymous):

it is a zero. 30 years. we just don't know r yet, so we have to find it

OpenStudy (anonymous):

first step is to divide, then take the log, then divide by 30 to get r

OpenStudy (anonymous):

ok just wanted to make sure. ok

OpenStudy (anonymous):

\[2.2358=e^{30t}\] \[\ln(2.2358)=30t\] \[t=\frac{\ln(2.2358)}{30}\]

OpenStudy (anonymous):

t=.0268 rounded steps clear so far?

OpenStudy (anonymous):

oh typos sorry. we are solving for r, not t !

OpenStudy (anonymous):

r = .0268 rounded. we are finding the rate

OpenStudy (anonymous):

ok, and I got .0268199702, so then as my final answer i got 23668677.39

OpenStudy (anonymous):

\[2.2358=e^{30r}\] \[\ln(2.2358)=30r\] \[r=\frac{\ln(2.2358)}{30}\] \[r=.0268\]

OpenStudy (anonymous):

not sure what you mean final answer

OpenStudy (anonymous):

the answer to a) is \[P(t)=10,586,223 e^{.0268 t}\]

OpenStudy (anonymous):

the answer to b we still have to compute.

OpenStudy (anonymous):

the answer to c) is what you get when you replace t by 53

OpenStudy (anonymous):

i mean the answer that i get when i plug in .0268 to \[10586223e^(30)(.0268)\]

OpenStudy (anonymous):

yes but you know that already. you already have you population in 1980

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

for part c just compute \[10,586,223e^{53\times .0268}\] and see what you get

OpenStudy (anonymous):

43860399.85

OpenStudy (anonymous):

for part b) you want doubling time. put \[2=e^{.0268t}\] and solve for t

OpenStudy (anonymous):

you get \[\ln(2)=.0268t\] \[t=\frac{\ln(2)}{.0268}\]

OpenStudy (anonymous):

or 25.86 so doubles about every 26 years or so

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

yw